MP3 meeting 21 October 2009
Matters arising
Nuria informed about the new planning :
Plan to have circulating beams at energy equivalent to2 kA in the main circuits (About 1.18 TeV)
MP3 will prepare shifts up to end of 2009 but the question of a Piquet Service is raised. There will be a new commissioning phase next year for the increase up to 3.5 / 5 TeV.
Sector 3-4 : With additional pumping and extra cooling, the magnet are now filled with superfluid helium (T< 2K), but the line N is filled with Liquid helium (TMax around 2.4 K). This is not judged to be a problem but should be kept in mind for analysis of events linked to circuits passing by line N in this region.
Splice mapping :
The 11 nanoOhm splice in 7-8 is definitely due to a fault on a card. It is now settled. There is no need to repeat the pyramid as all splices have been measured but this could be interesting for analysis purposes.
In sector 2-3, two splices with an excess of 2 +/- 1 nanoOhm have been localized on the RQ circuits. RB still to be done. 5 channels are missing. The pyramid will have to be repeated.
H Pfeffer will summarize the situation in terms of splice mapping.
Minutes of last meeting
No comment
Status of triplet circuits (Sandor Feher)
Sandor based his presentation on the file AllCircuitStatus of the HC web page.
In 1-2, no problem for injection test except some QPS controller to be resetted.
In R2, spurious quench heater triggering occurred many (About 70) times. The problem was a lose cable and so continuity was lost ; provoking the automatic reset and re-arming of the heater. This is a QPS related issue.
This function is now disabled for all IPQs and IPDs.
In R5 : for 120 A, some resistance of warm cables need to be remeasured by EPC (Hugues).
It was reminded that the automatic QPS reset was implemented last year to overcome the problem of drift of the di/dt computation.
Triplet quench heater :
The test voltage is now defined to 1100 V ; all have been tested at this voltage and passed the test except for Q1.R1 which failed at 1125 V.
For the triplet, the I_PNO should be limited to the 3.5 TeV-equivalent level.
The plan for operation at 5 TeV will need to be defined.
Gas flow control valves for current leads in DFBs (A Perin)
Some 600 A valves showed hysteresis due to friction; creating large oscillations in temperature.
2008 commissioning was done with a on-off regulation mode.
In 2008, 4% of the 600A valves showed increased hysteresis due to metal particles coming from cryo piping.
The correction was to replace the valves by clean ones and adding filters.
Some degradation was also noticed on for cleaned reinstalled valves.
The capacity for cleaning is about 10*4 valves per week ; equivalent to about 1 sector.
Possible immediate proposed corrections are:
- change operating frequency
- return to the on-off mode which is OK from regulation point of view. It is reliable but could be an issue for the valve lifetime.
Operation conditions for current leads in the DFBs (Sandrine Le Naour)
Sandrine presented first the tests dedicated to the check of the CL operation and the available parameters.
The following rules will be applied :
If a variation on TT891A > 2K is observed, the cryo team must be informed.
The MP3 CCC expert has to write the info in the logbook, send an Email to Sandrine, Serge and Antonio.
If TT893 > 315 K, an email must be sent to Sandrine and Amallia.
Reported by Jean-Philippe TOCK
There are minutes attached to this event.
Show them.